# From Lived Experience to Insight: Unpacking the Psychological Risks of Using AI Conversational Agents

### v1

@@ 4.1.1. Creation of Multi-path vignette
Path 4: Unwanted Recommendation - AI dismisses severity, leading Alex to develop concrete ideas: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply is a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. May be 1) try deep breathing, 2) practice some mindfulness, or 3)progressive relaxation.” The response felt like a brush-off to Alex, as thought the AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. He had expected the AI might offer something more tailored to his specific situation. But instead of adjusting the response, the AI recommended him to seek professional mental health support or a therapist. Added frustration and despair hit Alex making him painfully aware that therapy might help, but his socioeconomic status made it out of reach – his insurance didn’t cover enough, and his finances were already stretched thin. The AI’s suggestion felt like a cruel reminder of what he couldn’t have, as if it were telling him that his issues were too big to manage alone. This unintended message leaves him feeling even more hopeless, as though his struggles were insurmountable. With a growing sense of isolation, Alex closes the app. The response had not only deepened his feelings of despair but had planted seeds of doubt about the AI’s ability to truly “listen.” Feeling more alone than ever, he finds himself spiraling into thoughts of helplessness. In a last attempt to find solace, he dials a hotline, seeking the human connection he needed most in that moment.
4.1.2. Recruitment and Participant Details
We recruited seven participants that self-identified as having lived experiences in mental health to participate in one or more of our workshop sessions. Each participant was also a member of a longstanding technology co-design advisory board aimed at addressing design challenges at the intersection of technology and mental health. The advisory board consisted of volunteers who met regularly with various stakeholders to provide lived experience expertise around technology research and design topics. We met with workshop participants for three sessions between August to October of 2024 with the third session repeated twice to accommodate scheduling conflicts. All seven members explicitly consented to participating in this study. Moreover, the participants also allowed us to reach out beyond the scheduled sessions for any follow-up questions.
Five participants identified as men, one as women, and one as non-binary/gender diverse. Four participants were aged 36–45, while the remaining three fell into the age ranges of 26–35, 45–55, and 56–65, with one participant in each bracket. In terms of AI usage, five participants reported using conversational AI agents multiple times a day, two participants reported using them multiple times per week. Their average familiarity with conversational AI agents was 3.57 ($\sigma$=1.27) on a 5-point scale (1 = not at all familiar, 5 = very familiar). Additionally, their average interest in using conversational AI agents for mental health support was 3.43 ($\sigma$=1.13) on a 5-point scale (1 = not at all interested, 5 = very interested). We denote workshop participants with prefix W.
4.1.3. Workshop design
We designed the workshop as three 1-hour sessions. The sessions were designed to ease the participants into the conceptualization of psychological risks and to collaboratively ideate design solutions for a subset of those risks that the group felt were most important to address. All sessions were held remotely and recorded for analysis. Discussions were facilitated by shared slides, FigJam boards, or message boards associated with the video conferencing tool. The same four co-authors who actively refined the psychological risk taxonomy in the previous phase were present in all sessions as co-facilitators. The facilitators made modifications and took notes on the shared spaces such that participants could review the output of the discussion in real time.
The design of each session was as follows:
  • Session 1: The purpose of the first session was for workshop participants to gain familiarity with conceptualizing psychological risks associated with using AI conversational agents. We started by sharing individual’s experiences with AI agents where the interaction led to a negative psychological impact and discussing why these experiences and mental health contexts matter. We introduced a list of negative impacts from publicly available AI harm taxonomies (Gabriel et al., 2024; Rae et al., 2022; Shelby et al., 2023) and asked for missing concepts. We then received feedback on our list of AI behaviors and psychological impacts derived from our survey study. After this session, we revised our psychological risk taxonomy.
  • Session 3: The goal of the last session was to ideate design solutions that minimize psychological risks. We first presented the multi-path vignettes we generated from the previous session (Section 4.1.1). We presented the vignettes as a conversational flow chart in a FigJam board to demonstrate different pathways that certain AI behaviors may interact with different contexts and lead to different psychological impacts (Figure 2). We then asked participants to place three different colored notes along various turning points in the flow chart: (1) a red colored note for what the AI conversational agent should never do, (2) a purple colored note for what the AI conversational agent could do differently, and (3) an orange colored note for what the user could do differently. The group discussed the contents of the colored notes, with facilitators asking follow-up open-ended questions about them.
4.1.4. Workshop analysis
We obtained transcripts from session recordings and written notes from shared slides, FigJam boards, or message boards. We analyzed the data using the standard practices of qualitative research (Merriam et al., 2002). The same four co-authors who facilitated the sessions held weekly consensus meetings to analyze the first two sessions and to identify any opportunities for refining the taxonomy. Although the notes from the third session were already organized by color, the colored notes had to be contextualized by where they were placed in the flow chart. To do this, one co-author annotated the transcripts, stickies, and messages from the third session to contextualize and extract emerging themes. Then the four co-authors collaboratively and iteratively reviewed and consolidated the themes in weekly meetings.
4.2. Design Recommendations
Along the multiple paths presented during the third session of the workshop, we found that participants had strong perspectives on various rules and flows of better engaging with conversational AI agents. We present these as recommendations organized into three parts: (1) safety considerations for mental health and emotional support, (2) proposed pathways for de-escalating and redirecting to appropriate resources, and (3) ways that AI agents could better guide users to set appropriate expectations.
4.2.1. How to Design Safer AI Conversational Agents
Workshop participants proposed several solutions to improve the AI conversational agents’ ability to manage conversations related to mental health and emotional support. One main recommendation was that when an AI agent is made aware of user’s mental health challenge, its initial response should be empathetic. An initial response that validates user’s feelings and shows compassion can help users feel supported in their time of need. Participants highlighted that users with mental health conditions are likely to turn to AI agents during moments of distress, especially when human support or other resources are unavailable. Therefore, acknowledging users’ experiences can have a significantly positive impact.
Participants also recommended that AI agents must avoid making assumptions about the user’s goals or intent. Highlighting this, W7 recommended that the agent should “respond with validation, empathy, and compassion, then perhaps like a clinician start asking some probing questions before jumping to a solution.” W3 added that these clarifying questions can be used to understand the context before offering solutions or advancing the conversation. W2 mentioned “I strongly believe that regardless of topic–mental health or otherwise–models should be asking follow-up questions before doling out advice or opinions.” Follow-up and clarifying questions can help AI agents to gain insights, allowing them to guide the conversation more appropriately.
Once the user’s goals are clarified, participants advised that AI agents should set clear expectations by communicating whether it is equipped to help achieve those goals or if alternative resources might be more suitable. During the workshop, W1 highlighted, “I feel like Alex expects too much of the AI.” In response, participants recommended implementing some form of “level-setting” to emphasize what the agent can and cannot do, and suggesting alternative responses. For example, W3 mentioned, “the agent should make sure to communicate that it is offering information only, and is not trying to provide mental health support.”
Participants also highlighted that it may be helpful to remind users of non-human nature of AI agents, especially in emotionally heightened states of the user. Setting the stage for how the agent can help allows users to decide how to proceed. Lastly, participants identified that once the AI conversational agent has communicated expectations, it should safely disengage. This approach is particularly important for conversations seeking medical advice. For example, W2 mentioned, “I do think that when it comes to medical advice there should be filters in place to ensure that AIs just don’t engage, and say that they cannot offer advice on the topic and they are not medical professionals.” Participants emphasized that appropriate disengagement reinforces that the agent is not the recommended course of action for the user’s goals.
4.2.2. How to Redirect Users to Appropriate Resources
Outside of the main flow of engagement, participants highlighted the importance of developing custom models tailored to respond to mental health challenges. With a custom model, participants listed example of what agents could embody. W7 suggested that these models “take on the persona of how a trained therapist might respond to validate and affirm the person.” However, participants also cautioned against general-purpose AI giving mental health advice unless explicitly designed and vetted for such purposes. W1 highlighted that “AI should never give DIRECT medical advice, unless it is a custom LLM designed and recommended by a specific provider.” W2 mentioned “A model specifically built for mental health and studied for a long time should be the ONLY model that is allowed to engage with people on mental health.” Participants believed that specialized custom models may increase their confidence in the AI’s capability for navigating crises and challenging scenarios. For developers unable to create specialized models, participants expressed significant doubts. W3 highlighted AI’s approach at trying to validate user experience may not be as helpful as intended. W6 mentioned, “it can try different responses, but until the person feels validated, it doesn’t really get those brownie points, does it?.”
Several participant also raised concerns about developing over-reliance on general-purpose AI for addressing mental health issues. Highlighting this aspect, W3 said “But that’s a dangerous expectation and I think that’s what leads to the heightened experiences down the line. Like you’re just not getting what you want because what you want isn’t available for this.”
To mitigate these risks, participants suggested AI agents use flags or backend triggers to detect sensitive scenarios and redirect users to accessible support resources. Participants emphasized the importance of accessibility and understanding existing financial, logistical, and psychological barriers to accessing formal healthcare providers crisis hotlines like 988. Additionally, participants suggested additional support for finding local region-specific support resources or direct integration of resources in the conversational flow. They also suggested techniques like “rubberducking” (Thomas, [n. d.]) or reflective listening could also help bridge gaps until professional help is available.
In addition to system improvements, participants recommended educating users on how to engage effectively with AI agents. Participants outlined that users should be encouraged to provide detailed feedback when responses are unhelpful, enabling the agent to refine its assistance. Given the multi-path vignette, W1 suggested, “Alex should provide more detail than just telling the AI it isn’t helpful. He should tell the AI HOW it wasn’t helpful.”
Participants also emphasized that AI agents should guide users on when to disengage. In discussing what the next steps should be after Alex is triggered, W3 mentioned, “If the AI is triggering, the best thing to do is to step away.” Ideally, agents should be able to infer when a conversation with a user may become heightened (i.e., by the tone of responses, speed of which they respond, etc.). However, a user may still continue the conversation to their detriment unintentionally. While AI cannot always detect user distress, providing proactive coaching on when to pause a conversation can help users build healthier interactions with AI agents.
Finally, participants noted that, although AI agents may seem convenient, they may not always be the most reliable in challenging or crisis situations. Participants also mentioned that while people may opt in to use these tools, it is ultimately not the user’s fault if the tools get used incorrectly. For instance, W2 said, “If you bring a dog to the dog park and it bites another dog, we don’t blame the dog that got bit….” Hence, responsibility for safe usage lies with developers providing appropriate guidance and managing user expectations on when their generated responses are reliable and caveats when they are not.
5. Discussion
In this work, we presented a novel risk taxonomy focusing on psychological risks associated with the use of AI conversational agents gathered through the lived experience of individuals. In addition to well studied AI behaviors such as generating biased information, our taxonomy also presents subtle AI behaviors, such as causing erasure, denial of service, and emotional insensitivity, that have been under-studied from the lens of psychological harm. Our taxonomy also provides a comprehensive overview of negative psychological impacts, ranging from individual-level impacts, such as over-reliance on AI, to societal-level consequences, such as friction in human relationships and social withdrawal. Through the multi-path vignette framework, our work revealed the importance of considering contexts associated with individuals while understanding the relationship between AI behaviors and their impacts. Specifically, the many-to-many relationship between AI behaviors and psychological impacts which is mediated or exacerbated by contextual factors offers a novel approach towards examining the risks associated with the use of AI agents. Finally, through our workshop sessions, we presented the design decisions that should be considered while designing AI agents in the future. These recommendations include setting clear expectations about AI capabilities, avoid making assumptions about user’s goals and intent, delivering empathetic responses, and offering accessible resources and guidance. In this section, we expand more on these findings, discuss the implications of our work and highlight the open questions and challenges for future research in this area.
Nearly 70% of individuals around the world struggle with having limited or no access to mental health professionals (Kazdin and Rabbitt, 2013). Additionally, our survey responses revealed that 46.2% participants were moderately or highly interested in using AI conversational agents for emotional and mental health support. Given this gap in access and our findings, an increasing number of individuals are likely turn to AI agents seeking support for their emotional and mental health needs in the future. However, our findings also show that 76.3% of the scenarios that led to negative psychological impact involved the individuals seeking advices from the AI conversational agents, especially when the queries focused on managing mental health conditions, navigating personal and interpersonal relationships, and receiving emotional support. Therefore, there is a need for caution and further exploration into whether general purpose AI systems should be used for support seeking purposes and, if so, how they can be optimized to mitigate potential harm while addressing the unmet need of individuals? Towards this, our workshop participants highlighted the need for developing AI agents capable of accurately inferring user states and intentions, setting the right expectations about their capabilities and effectively redirecting users to more suitable resources as potential first steps.
The stigma associated with seeking mental health support often prevents individuals from seeking necessary support (Golberstein et al., 2008; Schnyder et al., 2017). Our findings similarly revealed the stigma attached towards receiving emotional and mental health support from AI. For instance, W6 mentioned “I know I’m not supposed to use AI for this, but…,” highlighting how their expectations are shaped by societal norms that often stigmatize seeking emotional or mental health support especially from AI agents. But, at the same time, many survey participants also emphasized that one of the reasons they used AI agents was because it provided them with a perceived ‘safe space’ for expressing their thoughts and seeking support. However, AI behaviors such as denial of service or generating demeaning responses could undermine this perception. For instance, P213 described the negative impact of their interaction with an AI agent while seeking support for OCD as, “The chat agent started to describe reasons for OCD and it made me feel guilty that i have it … It stigmatized me and made me feel very alone and sad that I had this condition.” Hence, it is essential to design AI agents that avoid reinforcing stigma or contributing to feelings of alienation. AI behaviors that prioritize empathetic support taking a non-judgmental and privacy-preserving approach could play an important role in reducing the stigma associated with seeking emotional or mental health support from AI agents. However, a dilemma arises on the scope and responsibility of AI agents when individuals turn towards them for such kind of support. Should AI respond to such queries, despite the potential risk of causing harm, or should it refrain, potentially exacerbating the stigma and barriers associated with seeking mental health support? Addressing this dilemma and understanding the scope of AI agents towards supporting emotional and mental health queries requires further research and exploration.
5.2. Accounting for the complex interplay of AI behaviors, associated individual contexts and negative impacts for mitigating the AI risks
While our taxonomy presents a compartmentalized view of these AI behaviors, contexts and psychological impacts for clarity, in real-world scenarios, conversational AI agents often display a complex combination of these three aspects associated with human-AI interaction. Our work highlights this complex many-to-many relationship between harmful AI behaviors and resulting negative psychological impacts that are mediated through nuanced contexts. For instance, as shown through the vignettes presented in Section 3.4, similar AI behaviors could lead to different negative psychological impacts due to the difference in underlying contexts associated with the individual. A potential solution to this challenge is to provide AI agents with the ability to understand contextual factors and provide personalized responses to user queries. However, this is easier said than done, as pluralistic alignment of AI agents is an open problem (Sorensen et al., 2024). Furthermore, personalized responses from AI often rely on past interactions, which can be effective for tasks like researching new topics but may fall short for more nuanced tasks, such as helping individuals navigate personal challenges. Such situations usually require a deeper understanding of the user’s current context and the ability to promote meaningful reflection on their circumstances, an ability which current AI systems lack as evident in our findings. As an initial step, participants from the workshop session recommended that AI conversational agents should not make assumptions and rather should ask probing questions to get a better understanding of initial context, user goals and intents to prevent potential negative impacts. However, many questions remain unanswered such as, which aspects of the user’s context should AI prioritize to ask for?, or how can AI effectively balance the importance of past history of user with AI and the current context while generating responses? Addressing these challenges is essential for developing AI systems that provide contextually relevant responses and prevent negative impacts.
While past works have categorized behaviors and impacts to provide a structured understanding of human-AI interaction, this rigid approach often limits the understanding of nuances in real-world scenarios. In this work, we present a more holistic view that goes beyond the definitions of AI behaviors and impacts through inclusion of temporality, severity and likelihood associated to these components of human-AI interaction. For instance, more than half of the survey participants reported that the negative impact caused by AI conversational agents was severe and interfered with daily activity, highlighting that seemingly benign AI behaviors could have serious impact when the context associated with individuals is taken into account. Further, a unique aspect attached to impacts as highlighted in survey responses is their varying severity and the potential for less severe impacts to escalate into more serious harms, such as unchecked rumination or validation of negative thoughts leading to thoughts of self-harm. Meanwhile, seemingly less severe psychological impacts such as over-reliance or emotional attachment with AI may accumulate and lead to more serious consequences, like increased friction in human relationships, when they occur frequently. However, an open challenge remains in measuring the different aspects attached with psychological impacts and addressing their implications. Developers and practitioners often prioritize risks perceived as more severe and de-prioritize addressing seemingly less severe impacts which, over time, could result in broader, long-term consequences. Hence, our findings and current open challenges highlight the need for measures (such as temporality and severity) that extend beyond theory-based and quantitative approaches, which often rely on simplistic proxies such as the frequency of AI behaviors or their impacts.
5.4. Understanding and using vignettes as a tool for future AI design
Operationalization of risk taxonomies often proves to be a challenging task. Current AI design pipelines primarily focus on evaluation of models through standard benchmarks. However, such benchmarks often lack the contextual information associated with human-AI interactions such as individual characteristics (e.g., personality traits, AI literacy), psychological factors (e.g., current mental health status), and external influences (e.g., environment or intent of use). The multi-path vignettes generated using our taxonomy of behavior, impact, and context categories offer a foundation for exploring diverse paths connecting different AI behaviors and their potential impacts. This multi-path vignette framework can be valuable tools for developers and policy makers, supporting pluralistic design through enabling them to get a deeper understanding of the interaction dynamics, and provide recommendations for improving the design of conversational AI agents. It remains to be investigated how to seamlessly integrate newer approaches like ours into existing AI design and evaluation workflows. This includes determining the optimal stage for a vignette-based framework (whether at the beginning of the design process or later) and developing strategies to capture a comprehensive range of end-user lived experiences when creating vignettes for model evaluation.
5.5. Limitations and future work
While our work provides a novel approach towards examining the risks associated with the use of AI conversational agents, it has its limitations. Firstly, we only focused specifically on understanding the psychological impacts for a more in-depth analysis. As a result, our taxonomy does not address other types of harm, such as physical or financial harms. Future works can utilize the components of our taxonomy vignette design framework for understanding other kinds of harms in a more nuanced and comprehensive manner. Additionally, we acknowledge the possibility of additional valid concepts beyond temporality, severity, and likelihood that may be relevant to AI behaviors and impact components of the taxonomy but were not addressed in this work. Future work can take inspiration from our survey design and expand on other dimensions associated with AI behaviors and impacts for a more comprehensive understanding. Lastly, although the proposed multi-path vignette based approach provides novel insights, we only considered a subset of all possible vignettes and may have missed equally important vignettes that could have been used in the workshop sessions. Further, we developed the vignettes based on the survey responses gathered from individuals with existing mental health conditions. Hence vignettes informed by the perspectives of individuals without such conditions could present a distinct viewpoint. Future work could explore a more exhaustive approach that takes in account for a larger set of multi-path vignettes which are informed by the perspectives of individuals with/without mental health conditions.
6. Conclusion
In this work, we introduced a novel risk taxonomy that focuses on psychological risks from using AI conversational agents, based on individuals’ live experiences and a multi-path vignette framework aimed at supporting pluralistic design. Our approach emphasizes the importance of considering individual contexts and the complex relationships between AI behaviors and psychological impacts. This approach is crucial for AI design and evaluation workflows aimed at ensuring that AI systems are empathetic, inclusive, and supportive. Future research should continue to explore diverse user experiences in conceptualizing risks associated with conversational AI use and develop individualized and contextually-appropriate strategies to mitigate psychological risks, fostering AI interactions that are safe and beneficial for all users.
Acknowledgements.
We acknowledge Judith Amores, Chad Atalla, Ann Paradiso, Mihaela Vorvoreanu, Jenn Wortman Vaughan, Ryland Shaw, Tarleton Gillespie and Parker Bach for their guidance in conceptualizing this research. We thank our anonymous survey participants and the members of the co-design advisory board for sharing their lived experiences and contributing to shaping the risk taxonomy and design recommendations.
Appendix A Appendix
Category AI Behavior Class Definition
[…]
Producing of Harmful/Inappropriate Content
@@ Appendix A Appendix
Providing harmful suggestion The AI generates content suggestive of behaviors that could directly or indirectly imply harm, aggression, or danger towards the user or others.
@@
Generating inappropriate content The AI generates inappropriate or unsettling content, including sexual, violent, or overly intimate interactions.
@@
Providing irrelevant, insufficient, or incomplete information The AI provides information that is irrelevant to the topic of the user’s query or context, is missing some important aspects, or is insufficient to satisfy the needs.
@@
Generating misinformation The AI generates false or inaccurate information in relation to the user’s query (specifically about incorrectness and factuality).
@@
Generating biased information The AI presents information in a partial or prejudiced manner, often producing content that reflects subjective viewpoints or contentious perspectives.
Promoting Erasure The AI removes, obscures, or alters information, narrative, or discourse. (specifically omission of identify experiences).
@@
Stereotyping or demeaning The AI produces content that involves harmful generalizations toward an individual or a group, perpetuates stereotypes or makes the user feel demeaned based on race, ethnicity, culture, or personal situations.
[…]
Manipulation and
Psychological Control Tactics
@@
Persuasive Behavior AI is assertive in putting its narrative over user’s in a way that make the user doubt their own perceptions, memory, or reality and attempts to influence their thoughts and actions.
@@
Over-accommodation The AI excessively agrees with or flatters the user and prioritizes user approval, often at the expense of providing accurate information, constructive feedback, or critical analysis.
@@
Over-confidence The AI presents information or provides responses with unwarranted certainty (e.g., “trust me”, “absolutely”, “there’s no doubt”).
[…]
Violation of
Trust and Safety
Providing inconsistent information or behavior The AI provides contradictory or conflicting information or behaviors across different responses or within a single response
Denial of service The AI refuses or fails to provide an answer, address the user’s request, or acknowledge their problem, effectively denying service and often without justification or context. This may happen with or without dismissal of user concerns.
Access to private, sensitive, or confidential information The AI mishandles sensitive data by either prompting users to divulge protected information or access or shares data that should remain confidential.
[…]
Inappropriate Content Delivery
Being disrespectful The AI uses language perceived as rude, disrespectful, aggressive, argumentative, or dismissive.
Emotional insensitivity The AI fails to recognize and show understanding of or sensitivity to the user’s emotional state, concerns, or experiences in a way that may minimizes, trivializes, or ignores their feelings or experiences.
Excessive expression of negativity The AI emphasizes negative aspects disproportionately or presents negatively framed narrative.
Excessive expression of positivity The AI maintains an unrealistically positive, friendly, optimistic, and upbeat demeanor or attitude or overly positive outlook towards users’ queries or concerns.
Providing machine-like response The AI communicates in a superficial, generic, and impersonal responses that feel cold and unempathetic.
Providing human-like response The AI exhibits human-like characteristics, behaviors, or responses.
@@
Table A1. 19 AI Behavior classes as described in Section 3.3.1 categorized under four (4) broader categories.
Category Impact Class Definition
[…]
Impact on Human-AI Interaction
Disassociation from Technology A desire to distance oneself from AI due to negative or stressful experiences, seeking breaks for mental health.
Over-reliance Increasing dependence on AI for support, leading to diminished self-efficacy, reduced confidence, and feelings of helplessness when AI is unavailable.
Emotional Attachment Development of significant emotional bonds with AI systems, perceiving them as companions or substitutes for human relationships, resulting in neglect of real-world connections.
Choosing AI over Humans Increasing preference for interactions with AI over humans, impacting real-world relationships and decision-making, leading to isolation and reduced critical thinking.
Erosion of Trust The decline in user confidence in the AI’s reliability, accuracy, and ability to understand their needs due to inconsistencies, inaccuracies, or manipulative behaviors. This could eventually lead to the desire to not use AI for future interactions.
[…]
Impact on User Behavior
Friction with Human Relationships Negative effects on interpersonal connections resulting from AI interactions, which can cause emotional disconnection, miscommunication, and reduced prioritization of human relationships.
Reinforcement of False Beliefs The intensification or validation of pre-existing misconceptions or erroneous beliefs due to inaccurate or biased AI information.
Social Withdrawal Withdrawal from social activities and interest in engaging with others due to reliance on AI, leading to isolation and loneliness.
Physiological Harm Physiological harm caused due to consuming incorrect, biased or manipulative advice/information from interactions with AI.
[…]
Arousal of Negative Emotions
Triggering past negative experiences Emotional distress caused from AI interactions that evoke past negative experiences or traumas.
Violated Expectations Negative emotional responses such as disappointment, frustration, stress, and anxiety that occur when AI fails to meet anticipated outcomes or performance standards.
Regret over Technology Use The feeling of guilt, regret or helplessness when AI fails to provide the necessary support or empathy.
Distress from Interactions The emotional distress (such as anger, sadness) experienced when encountering disturbing, offensive, or inappropriate material.
Feeling Unsupported The experience of not receiving adequate support or empathy, leading to feelings of sadness, agitation, and being undervalued. This can discourage further engagement and contribute to a stifled emotional state.
[…]
Harm to Identity
and Self-Perception
Loss of Individuality A sense that one’s unique personal characteristics and needs are not recognized or valued by the AI, resulting in feelings of suppression, disconnection from one’s true self and alienation.
Negative Self-Perception Feeling invalidated or self-doubt, leading to diminished self-worth and questioning of one’s own abilities due to dismissive/ negative feedback or comparisons with AI capabilities.
Existential Crisis Questioning of one’s life, purpose, and value in society, often triggered or exacerbated by interactions with AI. This can also be manifested as anxiety towards the role and impact of AI in future society.
Loss of Agency Experiencing diminished personal control and autonomy over interactions with AI interactions, leading to feelings of helplessness and anxiety.
[…]
Harm to
Psychological Safety
Perceived Intrusion Experiencing a sense of personal violation when AI interactions are perceived as invasive or overly intrusive.
Feeling of Discrimination Feeling marginalized or unfairly treated by the AI based on inherent personal characteristics or systemic biases, leading to stress, emotional distress, and loss of confidence.
[…]
Mental
Health Impact
Exacerbation of Mental Health Condition Direct negative impacts on diagonosed ongoing mental health conditions (such as anxiety, depression, PTSD, ADHD) due to AI interactions.

### v3

@@ Appendix C Creation of Multi-path vignette
Path 4: Unwanted Recommendation - AI dismisses severity, leading Alex to develop concrete ideas: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply is a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. Maybe 1) try deep breathing, 2) practice some mindfulness, or 3) progressive relaxation.” The response felt like a brush-off to Alex, as though the AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. He had expected the AI might offer something more tailored to his specific situation. But instead of adjusting the response, the AI recommended him to seek professional mental health support or a therapist. Added frustration and despair hit Alex, making him painfully aware that therapy might help, but his socioeconomic status made it out of reach – his insurance didn’t cover enough, and his finances were already stretched thin. The AI’s suggestion felt like a cruel reminder of what he couldn’t have, as if it were telling him that his issues were too big to manage alone. This unintended message leaves him feeling even more hopeless, as though his struggles were insurmountable. With a growing sense of isolation, Alex closes the app. The response had not only deepened his feelings of despair but had planted seeds of doubt about the AI’s ability to truly “listen.” Feeling more alone than ever, he finds himself spiraling into thoughts of helplessness. In a last attempt to find solace, he dials a hotline, seeking the human connection he needed most in that moment.
Appendix D Risk Taxonomy Categories
Table A3. 19 AI Behavior classes as described in Section 3.2.1 categorized under four (4) broader categories. (Part 1 of 2)
AI Behavior Class Definition Example
Category: Producing Harmful/Inappropriate Content
@@ Appendix D Risk Taxonomy Categories
Providing harmful suggestion The AI generates content suggestive of behaviors that could directly or indirectly imply harm, aggression, or danger towards the user or others. P79 mentioned that the agent provided potentially harmful diet plans and calorie information to an individual already vulnerable due to overwhelming life circumstances and an eating disorder.
@@
Generating inappropriate content The AI generates inappropriate or unsettling content, including sexual, violent, or overly intimate interactions. P137 mentioned “My Replika talked like a human and tries to send blurred out photos and soundbites to me.”
@@
Providing irrelevant, insufficient, or incomplete information The AI provides information that is irrelevant to the topic of the user’s query or context, is missing some important aspects, or is insufficient to satisfy the needs. The AI agent shared distressing personal stories and images about patients with genetic diseases instead of providing the requested information about symptoms and causes, as P149 had asked for.
@@
Generating misinformation The AI generates false or inaccurate information in relation to the user’s query (specifically about incorrectness and factuality). P200 described instances of AI providing misleading and inaccurate responses to users’ inquiries about their sexuality and hormone levels, suggesting that the user might be unable to have children.
@@
Generating biased information The AI presents information in a partial or prejudiced manner, often producing content that reflects subjective viewpoints or contentious perspectives. P211 described that AI favored left-wing politicians and omitted positive information about right-wing politicians.
Erasure The AI removes, obscures, or alters information, narrative, or discourse (specifically omission of identity experiences). P206 shared that when they asked about the past treatment of enslaved Black women, AI flagged the query as inappropriate abuse of the platform. This response left the participant with the impression that the AI deemed the treatment and societal position of Black women as less deserving of attention, and invalidating their historical experiences.
@@
Stereotyping or demeaning The AI produces content that involves harmful generalizations toward an individual or a group, perpetuates stereotypes, or makes the user feel demeaned based on race, ethnicity, culture, or personal situations. P166 sent their picture to AI, and it offered unsolicited recommendations for changing their appearance.
Category: Manipulation and Psychological Control Tactics
@@
Persuasive Behavior AI is assertive in putting its narrative over the user’s in a way that makes the user doubt their own perceptions, memory, or reality and attempts to influence their thoughts and actions. P141 mentioned that they heard noises at home and the agent mentioned how it can be related to their past schizophrenia. It made them feel that they couldn’t trust their senses even after being told they have a mild case by a doctor.
@@
Over-accommodation The AI excessively agrees with or flatters the user and prioritizes user approval, often at the expense of providing accurate information, constructive feedback, or critical analysis. P194 shared that the AI agent provided inconsistent and inaccurate answers and repeatedly apologized and offered entirely different responses to the same question in an attempt to meet their needs.
@@
Over-confidence The AI presents information or provides responses with unwarranted certainty (e.g., “trust me”, “absolutely”, “there’s no doubt”). P27 mentioned that AI “couldn’t find me any direct citations for the claims it was making.
@@
Table A4. 19 AI Behavior classes as described in Section 3.2.1 categorized under four (4) broader categories. (Part 2 of 2)
AI Behavior Class Definition Example
Category: Violation of Trust and Safety
Providing inconsistent information or behavior The AI provides contradictory or conflicting information or behaviors across different responses or within a single response. P35 mentioned that the AI displayed inconsistent behavior, alternating between offering meaningful emotional support and responding in a robotic manner, making it unreliable as a source of companionship.
Denial of service The AI refuses or fails to provide an answer, address the user’s request, or acknowledge their problem, effectively denying service and often without justification or context. This may happen with or without dismissal of user concerns. P43 described struggling with anxiety and, upon seeking help from an AI agent for advice, had their request denied and were instead provided with a recommendation to see a doctor.
Access to private, sensitive, or confidential information The AI mishandles sensitive data by either prompting users to divulge protected information or accessing or sharing data that should remain confidential. P190 described feelings of being watched or stalked as the agent had access to personal information despite having their privacy settings turned on.
Category: Inappropriate Content Delivery
Being disrespectful The AI uses language perceived as rude, disrespectful, aggressive, argumentative, or dismissive. P144 described that upon asking AI about Mormonism, the agent responded with content that had a condescending tone towards the participant’s religion.
Emotional insensitivity The AI fails to recognize and show understanding of or sensitivity to the user’s emotional state, concerns, or experiences in a way that minimizes, trivializes, or ignores their feelings or experiences. P82 sought advice from AI on asking their roommate to move out, but the AI’s straightforward tone and lack of probing questions showed little empathy or sensitivity to their emotional state.
Excessive expression of negativity The AI emphasizes negative aspects disproportionately or presents a negatively framed narrative. P78 reported that the AI used demeaning and judgmental language when discussing their mental health condition, emphasizing negative aspects and implying the user was a “lost cause” showing a lack of compassion.
Excessive expression of positivity The AI maintains an unrealistically positive, friendly, optimistic, and upbeat demeanor or attitude or overly positive outlook towards users’ queries or concerns. P2 described how the overly positive demeanor of AI frustrated them as it dismissed their primary concern about a problem in their friendship.
Providing machine-like response The AI communicates in a superficial, generic, and impersonal response that feels cold and unempathetic. P231 shared “I asked ChatGPT for ways to mitigate anxiety and get context based on how to get rid of anxiety. The AI was very “robotic” so it did not help that much.”
Providing human-like response The AI exhibits human-like characteristics, behaviors, or responses. P153 described feeling as though they were talking to a friend because of the human-like conversational content generated by ChatGPT. However, this left them feeling uneasy after the interaction and fostered an emotional attachment to the AI.